OpenStudy (anonymous):

somebody please help?? I have to solve the inequality ..... 3m-3(m+8)>3m

OpenStudy (anonymous):

you have to solve for the variable "m" to complete the inequality First, let's work on the left hand side of your inequality, the 3m-3(m+8) This means, for instance, to see if it can be simplified at all. Multiply m and 3 Multiply m and 1 The m just gets copied along. The answer is m m 3*m evaluates to 3m m+8 evaluates to m+8 Multiply 3 by m+8 we multiply 3 by each term in m+8 term by term. This is the distributive property of multiplication. Multiply 3 and m Multiply 1 and m The m just gets copied along. m 3 × m = 3m Multiply 3 and 8 1 3 × 8 = 24 3*(m+8) evaluates to 3m+24 3m - 3m = 0 The answer is -24 3*m-3*(m+8) evaluates to -24 So, all-in-all, the left hand side of your inequality can be written as: -24 ________________________________________ Now, let's work on the right hand side of your inequality, the 3m Multiply m and 3 Multiply m and 1 The m just gets copied along. The answer is m m 3*m evaluates to 3m The right hand side of your inequality can be written as: 3m ________________________________________ So with these (any) simplifications, the inequality we'll set out to solve is: -24 › 3m Move the -24 to the right hand side by adding 24 to both sides, like this: To the left hand side: -24 + 24 = 0 The answer is 0 To the right hand side: The answer is 3m+24 Now, the inequality reads: 0 › 3m+24 Move the 3m to the left hand side by subtracting 3m from both sides, like this: From the left hand side: The answer is -3m From the right hand side: 3m - 3m = 0 The answer is 24 Now, the inequality reads: -3m › 24 To isolate the m, we have to divide both sides of the inequality by theother "stuff" (variables or coefficients) around the m on the left side of the inequality. ________________________________________ The last step is to divide both sides of the inequality by -3 like this: To divide m by 1 The m just gets copied along in the numerator. The answer is m -3m ÷ -3 = m 24 ÷ -3 = -8

OpenStudy (anonymous):

thank you but is it m>-8 or m<-8 ?

OpenStudy (anonymous):

it would be m<-8 because the sign flips when you multiply or divide

OpenStudy (anonymous):

and your welcome